OpenStudy (faolan):

I WILL FAN N MEDAL!!! I JUST NEED HELP PLS?!?!?!?!!! Find the equation of the ellipse with the following properties. The ellipse with foci at (4, 0) and (-4, 0); y-intercepts (0, 3) and (0, -3).

OpenStudy (faolan):

@jigglypuff314 @mathmale KNow how to do this?

OpenStudy (faolan):

@bigpapa1 @random231 @AloneS

random231 (random231):

do you know thw general eqution of ellipse?

OpenStudy (faolan):

no.....

OpenStudy (4everaddicted2anime):

The general equation of ellipse is \[\frac{ x^2 }{ a^2 }+\frac{ y^2 }{ b^2 }=1\] x and y are the coordinates of any point on the ellipse. a and b are the radius on the x and y axes respectively

OpenStudy (faolan):

I got it thx. substitute teacher actually knew something for once